# Hetero-EUCLID: Interpretable model discovery for heterogeneous hyperelastic materials using stress-unsupervised learning11footnotemark: 1

### v1

@@ 3.2 Pattern case study
We then solve the system using a Bayesian framework with spike-slab priors, employing three parallel MCMC chains (Markov Chain Monte Carlo) of length 500 each (refer A and Joshi et al. [46] for details on the Bayesian framework). The outcome of this procedure is a segment-wise interpretable discovery of material models. Table 2 benchmarks the predicted model coefficients $\theta_{1}$ and $\theta_{6}$ against the ground truth values for the $\text{NH2}_{\text{a}}$$\text{NH2}_{\text{b}}$ material combination (Equation 29). It can be observed that, with only 2% of the domain nodal data sampled, the identified physically interpretable coefficients as shear and bulk moduli closely match their true values. For coefficients $\theta_{2}$ to $\theta_{5}$, which are not expected to contribute under the Neo-Hookean model formulation, the Bayesian framework predicts zero values due to the sparsity-promoting nature of the spike-slab prior. The small standard deviation values further indicate high confidence in the inferred parameters.
The influence of noise on Hetero-EUCLID for cross-shaped inclusion is reported in Figure 9. Under high noise case $(\sigma_{u,\text{high}}=5\times 10^{-3})$, the effect on the residual force norm $(f^{a}_{\text{res}})$ is shown in Figure 9e, which may be visually compared with the noise-free case in Figure 8e. Despite the added noise, the material region segmentation remains accurate as illustrated in Figure 9f. Supplementary Figure S1 shows the corresponding visualization results for moderate noise condition $(\sigma_{u,\text{mod}}=2\times 10^{-3})$. Table 3 reports the predicted material parameters for both moderate and high noise cases, where the maximum absolute error remains below 1.2% compared to the ground truth. Supplementary section S2 presents an empirical validation study by comparing the strain energy density along six different deformation paths between the discovered material model and the corresponding ground truth. Supplementary Figure S5 shows the comparative strain energy density plots for the cross-shaped inclusion pattern under high noise conditions. The results indicate good agreement between the predicted and true strain energy densities, demonstrating the robustness of the proposed method under noise conditions.
Type 2: Split-domain pattern
We further investigate a more complex heterogeneous configuration, featuring a split-domain pattern with three distinct material regions whose interfaces converge at a common junction located on one of the external boundaries, as illustrated in 7(b). The domain is discretized using 10,972 wedge elements, and the underlying ground truth material model comprises of a combination of Neo-Hookean, Isihara, and Haines–Wilson constitutive laws. Using the same non-equi-biaxial loading conditions as in the previous case, we perform inverse parameter identification with the Hetero-EUCLID framework under noise-free, high, and moderate noise conditions. The corresponding results are presented in Figure 10, 11, and Supplementary Figure S2, respectively. In each of the above scenarios, the residual force distribution $(f^{a}_{\text{res}})$ reveals clear mechanical imbalance at the material interfaces, as shown in Figure 10e, 11e and S2e. Next, the segmentation step is executed using the seed-based growth algorithm with a tunable threshold parameter $\lambda$ (refer to Equation 26), and the resulting predicted material segments are shown in Figure 10f, 11f, and S2f. In each case, a small degree of misclustering is observed at the junction of the material interfaces, affecting only a few elements.
Pattern Type: Split-domain $|$ Mesh elements (C3D6): 10,972 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth Refer (a) 1 NH2 6.00 - - - - 32.00 2 ISH 4.00 0.50 0.30 - - 21.00 - 3 HW 1.00 0.15 - 0.02 0.00 10.00 No Noise Refer (f) 3 NH2 6.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 32.00 $\pm$ 0.00 2 ISH 3.99 $\pm$ 0.03 0.51 $\pm$ 0.01 0.29 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 20.99 $\pm$ 0.01 2% 4 HW 0.93 $\pm$ 0.09 0.21 $\pm$ 0.02 0.00 $\pm$ 0.01 0.02 $\pm$ 0.00 0.00 $\pm$ 0.00 10.05 $\pm$ 0.01
Table 4: True vs. predicted material parameters for split-domain pattern without noise using Bayesian-EUCLID framework.
Figure 10: Model discovery for the split-domain pattern (7(b)), with sub-figures (b)–(g) corresponding to the zero noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.
@@
Pattern Type: Split-domain $|$ Mesh elements (C3D6): 10,972 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth Refer (a) 1 NH2 6.00 - - - - 32.00 2 ISH 4.00 0.50 0.30 - - 21.00 - 3 HW 1.00 0.15 - 0.02 0.00 10.00 Moderate Noise ($\sigma_{u}=2\times 10^{-3}$) 3 NH2 5.98 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 31.79 $\pm$ 0.03 2 ISH 5.25 $\pm$ 0.16 0.20 $\pm$ 0.05 0.21 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 21.89 $\pm$ 0.05 2% 4 HW 1.62 $\pm$ 0.05 0.28 $\pm$ 0.01 0.00 $\pm$ 0.01 0.02 $\pm$ 0.00 0.00 $\pm$ 0.00 9.72 $\pm$ 0.03 High Noise ($\sigma_{u}=5\times 10^{-3}$) Refer (f) 3 NH2 5.98 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 31.62 $\pm$ 0.05 2 ISH 6.14 $\pm$ 0.20 0.00 $\pm$ 0.01 0.07 $\pm$ 0.08 0.00 $\pm$ 0.00 0.01 $\pm$ 0.01 21.80 $\pm$ 0.08 2% 4 HW 1.70 $\pm$ 0.05 0.25 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 9.57 $\pm$ 0.04
@@
Table 5: True vs. predicted material parameters for split-domain pattern with noise using Bayesian-EUCLID framework.
@@
Figure 11: Model discovery for the split-domain pattern (7(b)), with sub-figures (b)–(g) corresponding to the high noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.
@@
Table 4 compares the predicted material parameters $\boldsymbol{\theta}$ under noise-free conditions against the ground truth values. The parameters corresponding to the Neo-Hookean (NH2: $\theta_{1,6}$), Isihara (ISH: $\theta_{1-3,6}$), and Haines–Wilson (HW: $\theta_{1,2,4-6}$) models are all recovered with high accuracy and low standard deviation. Under noisy conditions, as reported in Table 5, the region governed by the NH2 model is predicted accurately. However, the regions corresponding to ISH and HW models exhibit some deviations in the predicted parameters, likely due to the compounded effect of noise and minor misclustering. The noise-induced deviation can also be observed from the strain energy density plots for HW under high noise conditions, as shown in Supplementary Figure S6. The prediction accuracies can be expected to improve with the use of data from multiple experiments/load-steps instead of a single non-equi-biaxial stretch test.
Type 3: Multiple inner inclusions
@@
Type 4: Cahn-Hilliard pattern
In this type of domain heterogeneity, we investigate the Cahn-Hilliard pattern, which is often used to represent the biological pattern formation in soft tissue mechanics. 7(d) illustrates this configuration, featuring five adjacent material regions alternately composed of two Neo-Hookean ($\text{NH2}_{\text{a}},\text{NH2}_{\text{b}}$) constitutive models. This specimen serves to evaluate the ability of Hetero-EUCLID to distinguish material sub-regions when multiple zones share identical mechanical properties. The domain is meshed with 21,040 uniformly distributed wedge-type elements. Using ABAQUS/CAE, we generate the full-field computational displacement dataset for quasi-static non-equi-biaxial loading as shown in Figure 14b-d. The Cahn–Hilliard heterogeneous configuration is analyzed under noise-free, high, and moderate-noise conditions, with the corresponding results summarized in Figure 14, 15, and Supplementary Figure S4, respectively.
Pattern Type: Multiple inner inclusions $|$ Mesh elements (C3D6): 10,816 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth Refer (a) 1 ISH 4.00 0.50 0.30 - - 21.00 - 2 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 3 HW 1.00 0.15 - 0.02 0.00 10.00 4 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 No Noise Refer (f) 1 ISH 4.00 $\pm$ 0.00 0.50 $\pm$ 0.00 0.30 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 21.01 $\pm$ 0.00 10% 4 $\text{NH2}_{\text{a}}$ 5.40 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.99 $\pm$ 0.01 2 HW 0.72 $\pm$ 0.07 0.00 $\pm$ 0.03 0.00 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 9.81 $\pm$ 0.01 3 $\text{NH2}_{\text{b}}$ 1.75 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.83 $\pm$ 0.01
Table 6: True vs. predicted material parameters for multiple inner inclusions pattern without noise using Bayesian-EUCLID framework.

### v2

@@ 3.3 Pattern case study
We then solve the system using a Bayesian framework with spike-slab priors, employing three parallel MCMC chains (Markov Chain Monte Carlo) of length 500 each (refer A and Joshi et al. [2022] for details on the Bayesian framework). The outcome of this procedure is a segment-wise interpretable discovery of material models. Table 2 benchmarks the predicted model coefficients $\theta_{1}$ and $\theta_{19}$ against the ground truth values for the $\text{NH2}_{\text{a}}$$\text{NH2}_{\text{b}}$ material combination (Equation 29). It can be observed that, with only 2% of the domain nodal data sampled, the identified physically interpretable coefficients, namely shear and bulk moduli, closely match their true values. For coefficients $\theta_{2}$ to $\theta_{5}$, which are not expected to contribute under the Neo-Hookean model formulation, the Bayesian framework predicts zero values due to the sparsity-promoting nature of the spike-slab prior. The small standard deviation values further indicate high confidence in the inferred parameters.
@@
Pattern Type 1: Cross-shaped inclusion $|$ Mesh elements (C3D6): 10,442 $|$ Activated feature library terms, $\boldsymbol{n_{\mathit{f}}=6}$ Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{19}$ (MPa) Free nodes sub-sampling Ground Truth Refer 8(a) 1 $\text{NH2}_{\text{a}}$ 1.80 - - - - 6.00 - 2 $\text{NH2}_{\text{b}}$ 5.40 - - - - 15.00 No Noise Refer 8(f) 1 $\text{NH2}_{\text{a}}$ 1.79 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.00 $\pm$ 0.00 2% 2 $\text{NH2}_{\text{b}}$ 5.46 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.13 $\pm$ 0.04 Moderate Noise, S1(f) ($\sigma_{u}=2\times 10^{-3}$) 1 $\text{NH2}_{\text{a}}$ 1.81 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.01 $\pm$ 0.00 2% 2 $\text{NH2}_{\text{b}}$ 5.42 $\pm$ 0.07 0.01 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.92 $\pm$ 0.13 High Noise, 8(i) ($\sigma_{u}=5\times 10^{-3}$) 2 $\text{NH2}_{\text{a}}$ 1.81 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.98 $\pm$ 0.00 2% 3 $\text{NH2}_{\text{b}}$ 5.25 $\pm$ 0.09 0.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.40 $\pm$ 0.22
@@
Table 2: True vs. predicted material parameters for cross-shaped inclusion pattern (7(a)) across different noise levels using Bayesian-EUCLID framework.
The influence of noise on Hetero-EUCLID for cross-shaped inclusion is reported in Figure 8. Under high noise case $(\sigma_{u,\text{high}}=5\times 10^{-3})$, the effect on the residual force norm $(f^{a}_{\text{res}})$ is shown in Figure 8h, which may be visually compared with the noise-free case in Figure 8e. Despite the added noise, the material region segmentation remains accurate as illustrated in Figure 8i. Supplementary Figure S1 shows the corresponding visualization results for moderate noise condition $(\sigma_{u,\text{mod}}=2\times 10^{-3})$. Table 2 reports the predicted material parameters for both moderate and high noise cases, where the maximum absolute error remains below 4% compared to the ground truth. Supplementary section S2 presents an empirical validation study by comparing the strain energy density along six different deformation paths between the discovered material model and the corresponding ground truth. Supplementary Figure S7 shows the comparative strain energy density plots for the cross-shaped inclusion pattern under high noise conditions. The results indicate good agreement between the predicted and true strain energy densities, demonstrating the robustness of the proposed method under noise conditions. In G, we repeat the study for a Type-1 geometry (cross-shaped inclusion), with the matrix and inclusion materials having similar material properties to evaluate the ability of the Hetero-EUCLID framework to accurately characterize and segment materials with similar properties.
@@
Figure 8: Model discovery for the cross-shaped inclusion pattern (7(a)). The ground truth material segment IDs are shown in (a), and the displacement fields in the $x$, $y$, and $z$ directions are provided in (b)–(d). For the zero-noise case, (e) depicts the residual force norm distribution (Equation 26) from the homogenized model, (f) illustrates the material segments identified through the segmentation algorithm, and (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ set used to construct the heterogenized model. The high-noise case follows a similar structure in panels (h)–(j), showing how the residual force field, segmentation, and sub-sampling behave under high noise levels. In both scenarios, the final heterogenized model is solved using the Bayesian framework to recover the region-wise constitutive parameters as listed in Table 2.
Pattern Type 2: Split-domain pattern
We further investigate a more complex heterogeneous configuration, featuring a split-domain pattern with three distinct material regions whose interfaces converge at a common junction located on one of the external boundaries, as illustrated in 7(b). The domain is discretized using 10,972 wedge elements, and the underlying ground truth material model comprises a combination of Neo-Hookean, Isihara, and Haines–Wilson constitutive laws. Using the same non-equi-biaxial loading conditions as in the previous case, we perform inverse parameter identification with the Hetero-EUCLID framework under noise-free, high-noise, and moderate-noise conditions. The corresponding results are presented in Figure 9e-g, 9h-j and Supplementary Figure S2e-g, respectively. In each of the above scenarios, the residual force distribution $(f^{a}_{\text{res}})$ reveals a clear mechanical imbalance at the material interfaces, as shown in Figure 9e, 9h and S2e. Next, the segmentation step is executed using the seed-based growth algorithm with a tunable threshold parameter $\lambda$ (refer to Equation 28), and the resulting predicted material segments are shown in Figure 9f, 9i, and S2f. In each case, a small degree of misclustering is observed at the junction of the material interfaces, affecting only a few elements.
@@
Table 3 compares the predicted material parameters $\boldsymbol{\theta}$ under noise-free conditions against the ground truth values. The parameters corresponding to the Neo-Hookean (NH2: $\theta_{1,19}$), Isihara (ISH: $\theta_{1-3,19}$), and Haines–Wilson (HW: $\theta_{1,2,4,5,19}$) models are all recovered with high accuracy and low standard deviation. Under noisy conditions, as reported in Table 3, the region governed by the NH2 model is predicted accurately. However, the regions corresponding to ISH and HW models exhibit some deviations in the predicted parameters, likely due to the compounded effect of noise and minor misclustering. The noise-induced deviation can also be observed from the strain energy density plots for HW under high noise conditions, as shown in Supplementary Figure S8. The prediction accuracies can be expected to improve with the use of data from multiple experiments/load-steps instead of a single non-equi-biaxial stretch test.
Pattern Type 2: Split-domain $|$ Mesh elements (C3D6): 10,972 $|$ Activated feature library terms, $\boldsymbol{n_{\mathit{f}}=6}$ Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{19}$ (MPa) Free nodes sub-sampling Ground Truth Refer 9(a) 1 NH2 6.00 - - - - 32.00 2 ISH 4.00 0.50 0.30 - - 21.00 - 3 HW 1.00 0.15 - 0.02 0.00 10.00 No Noise ($\sigma_{u}=0$) Refer 9(f) 3 NH2 6.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 32.00 $\pm$ 0.00 2 ISH 3.99 $\pm$ 0.03 0.51 $\pm$ 0.01 0.29 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 20.99 $\pm$ 0.01 2% 4 HW 0.92 $\pm$ 0.12 0.22 $\pm$ 0.03 0.00 $\pm$ 0.02 0.02 $\pm$ 0.00 0.00 $\pm$ 0.00 10.05 $\pm$ 0.01 Moderate Noise ($\sigma_{u}=2\times 10^{-3}$) Refer S2(f) 3 NH2 5.99 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 31.81 $\pm$ 0.02 2 ISH 5.25 $\pm$ 0.10 0.21 $\pm$ 0.03 0.21 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 21.86 $\pm$ 0.03 2% 4 HW 1.59 $\pm$ 0.09 0.28 $\pm$ 0.01 0.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 9.71 $\pm$ 0.02 High Noise ($\sigma_{u}=5\times 10^{-3}$) Refer 9(i) 3 NH2 5.99 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 31.63 $\pm$ 0.04 2 ISH 6.12 $\pm$ 0.22 0.01 $\pm$ 0.02 0.07 $\pm$ 0.08 0.00 $\pm$ 0.00 0.01 $\pm$ 0.01 21.79 $\pm$ 0.07 2% 4 HW 1.69 $\pm$ 0.04 0.25 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 9.56 $\pm$ 0.04
Table 3: True vs. predicted material parameters for split-domain pattern (7(b)) across different noise levels using Bayesian-EUCLID framework.
Figure 9: Model discovery for the split-domain pattern (7(b)). The ground truth material segment IDs are shown in (a), and the displacement fields in the $x$, $y$, and $z$ directions are provided in (b)–(d). For the zero-noise case, (e) depicts the residual force norm distribution (Equation 26) from the homogenized model, (f) illustrates the material segments identified through the segmentation algorithm, and (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ set used to construct the heterogenized model. The high-noise case follows a similar structure in panels (h)–(j), showing how the residual force field, segmentation, and sub-sampling behave under high noise levels. In both scenarios, the final heterogenized model is solved using the Bayesian framework to recover the region-wise constitutive parameters as listed in Table 3.
Pattern Type 3: Multiple inner inclusions
@@
Pattern Type 3: Multiple inner inclusions $|$ Mesh elements (C3D6): 10,816 $|$ Activated feature library terms, $\boldsymbol{n_{\mathit{f}}=6}$ Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{19}$ (MPa) Free nodes sub-sampling Ground Truth Refer 10(a) 1 ISH 4.00 0.50 0.30 - - 21.00 - 2 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 3 HW 1.00 0.15 - 0.02 0.00 10.00 4 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 No Noise ($\sigma_{u}=0$) Refer 10(f) 1 ISH 4.00 $\pm$ 0.00 0.50 $\pm$ 0.00 0.30 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 21.00 $\pm$ 0.00 10% 4 $\text{NH2}_{\text{a}}$ 5.40 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.99 $\pm$ 0.02 2 HW 0.72 $\pm$ 0.01 0.22 $\pm$ 0.00 0.00 $\pm$ 0.00 0.03 $\pm$ 0.00 0.00 $\pm$ 0.00 9.81 $\pm$ 0.02 3 $\text{NH2}_{\text{b}}$ 1.75 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.83 $\pm$ 0.01 Moderate Noise ($\sigma_{u}=2\times 10^{-3}$) Refer S3(f) 2 ISH 4.10 $\pm$ 0.11 0.51 $\pm$ 0.00 0.29 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 21.12 $\pm$ 0.05 10% 5 $\text{NH2}_{\text{a}}$ 5.34 $\pm$ 0.12 0.00 $\pm$ 0.01 0.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.83 $\pm$ 0.22 3 HW 0.41 $\pm$ 6.61 0.52 $\pm$ 0.07 0.01 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 9.24 $\pm$ 9.14 4 $\text{NH2}_{\text{b}}$ 0.06 $\pm$ 0.24 0.08 $\pm$ 0.02 0.02 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 2.81 $\pm$ 0.42 High Noise ($\sigma_{u}=5\times 10^{-3}$) Refer 10(i) 2 ISH 5.16 $\pm$ 2.65 0.48 $\pm$ 0.10 0.17 $\pm$ 0.07 0.00 $\pm$ 0.01 0.00 $\pm$ 0.00 21.02 $\pm$ 0.27 10% 5 $\text{NH2}_{\text{a}}$ 4.91 $\pm$ 0.34 0.00 $\pm$ 0.03 0.00 $\pm$ 0.03 0.26 $\pm$ 6.29 0.00 $\pm$ 0.00 13.84 $\pm$ 0.92 3 HW 1.92 $\pm$ 7.30 0.62 $\pm$ 0.09 0.00 $\pm$ 0.04 0.03 $\pm$ 0.07 0.00 $\pm$ 0.00 6.15 $\pm$ 1.58 4 $\text{NH2}_{\text{b}}$ 0.67 $\pm$ 0.51 0.05 $\pm$ 0.05 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.13 $\pm$ 0.79
Table 4: True vs. predicted material parameters for multiple inner inclusions pattern (7(c)) across different noise levels using Bayesian-EUCLID framework.